OpenStudy (anonymous):

Write the point-slope form of the equation passing through (5, -1) with a slope of 6. Answer y-1=6(x+5) y+5=6(x-1) y+1=6(x-5) y-5=6(x+1)

OpenStudy (hba):

Use the formula, \[\huge\ y-y_1=m(x-x_1)\] Where, \[\huge\ (x_1,y_1)=(5,-1),m=6\]

OpenStudy (hba):

@mikaela19900630 This is gonna be some positive energy running between us if you reply ?

jimthompson5910 (jim_thompson5910):

Similar Problem: Write the point-slope form of the equation passing through (16, -5) with a slope of 3. m = 3 (x1, y1) = (16, -5) x1 = 16 y1 = -5 Equation: y - y1 = m(x - x1) y - (-5) = 3(x - 16) y + 5 = 3(x - 16)

jimthompson5910 (jim_thompson5910):

This is just an example, not the answer to your question (it's very similar though)

OpenStudy (anonymous):

okay so i just plug the #s in right @hba

OpenStudy (hba):

Yeah right :) Easy as that :)

OpenStudy (anonymous):

okay so were does the extra y and x come in? if i only have a slope and one set of coordinates ??

OpenStudy (hba):

Let the y and x be as they are.

OpenStudy (anonymous):

so y-1=6(x+5)

OpenStudy (hba):

No, See this (-a)(-b)=+ab

OpenStudy (hba):

@mikaela19900630

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

im confused now

OpenStudy (anonymous):

@hba

OpenStudy (hba):

(y-(-1) ) =6(x-5)

OpenStudy (hba):

@mikaela19900630 Is that clear now ?

OpenStudy (anonymous):

isnt that what i said!?!?!

OpenStudy (hba):

No you were confused with the signs :)

OpenStudy (anonymous):

oh okay i think i got it now thank you